OpenStudy (anonymous):

x+y=8 y-x+2 I need to find the slope and Y intercept, can someone please show me step by step how to solve this!

OpenStudy (anonymous):

these are two separate lines. the second one is missing an = sign

OpenStudy (anonymous):

do you need to find the slopes and y intercepts of these two lines?

OpenStudy (anonymous):

my mistake it is y-x=2

OpenStudy (anonymous):

okay. the first line can be written in the slope intercept form y = mx +b x+y = 8 subtracting x from both sides,we get y = -x+8 y = (-1)x +8 so the slope is -1 and the y intercept is 8 similarly do the other one.